Please Draw this and explain28. As a 2.0-kg block travels around a 0.50-m radius circle it has an angular speed of 12 rad/s. The circle is parallel to th
Answer #1

given

mass m = 2 kg

angular speed w = 12 rad/s

linear speed v = r*w = 0.5*12 = 6 m/s

from the relation

angular momentum L = mvr'

here r' = sqrt(0.75^2+0.5^2) = 0.901 m

L = 2*6*0.901 = 10.82 kgm^2/s

option c